Caroline can sketch 2 cartoon strips in five hours. How long will it take her to sketch 8 ​strips?

Found 2 solutions by josgarithmetic, Alan3354:
Answer by josgarithmetic(39629)

Four times as long.

Answer by Alan3354(69443)

I think some cartoon strips take more time than others, but, assuming the times per strip is a constant:
=============
4 times as long.
Not "4 times longer"
It the time for a strip is t:
1 time longer = t + t = 2t
2 times longer = t + 2t = 3t
3 times longer = t + 3t = 4t
4 times longer = t + 4t = 5t
--------------
Know what I'm saying?